OpenStudy (anonymous):

Which pair of ratios form a proportion? A. 5 : 11 and 6 : 12 B. 7 : 13 and 21 : 39 C. 8 : 14 and 24 : 30 D. 10 : 20 and 20 : 30

OpenStudy (phebe):

D im sure

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

thanks!!! can u help me with another one?

OpenStudy (phebe):

yess @sggaa

OpenStudy (phebe):

medal plzz

OpenStudy (anonymous):

How is it D?

OpenStudy (anonymous):

@sggaa The answer isn't D

OpenStudy (anonymous):

10/20 = 0.5, 20/30 = 0.66

OpenStudy (phebe):

ohh ok well my bad im wrong

OpenStudy (anonymous):

Yes. 7*3=21. 13*3=39. The answer would be B

OpenStudy (anonymous):

do 1/8 and 3/20 form a porportion and how do I give metals

OpenStudy (anonymous):

You click best response. Its the blue button

OpenStudy (anonymous):

And no. 1/8 and 3/20 isn't a proportion :)

OpenStudy (anonymous):

Yup, it's B.

OpenStudy (anonymous):

How do I determine whether a ratio forms a proprotion

OpenStudy (anonymous):

Just divide

OpenStudy (anonymous):